5 lollipops and 12 toffees have a mass of 61 grams
10 lollipops and 13 toffees have a mass of 89 grams
I. If the mass of one lollipop is x grams and the mass of one toffee is y grams, write 2 equations in x and y to represent the above information
ii. Calculate the mass of:

A) One lollipop

B) One toffee

Answer by Boreal(15235)

5x+12y=61
10x+13y=89
-10x-24y=-122, multiplying the first equation by -2
-11y=-33
y=3 gms for toffee
then x=5 gms, checking in equation 1
-
see if they work in the second equation.
They do
Lollipop is 5 gm
Toffee is 3 gm